# Tree-Wasserstein Distance for High Dimensional Data with a Latent Feature Hierarchy

### v2

@@ A.4 Preliminaries on Trees
Definition A.8 (LCA Relation [Wang and Wang, 2018].).
@@ Appendix B Additional Related Work
Wasserstein Distance on Hyperbolic Manifold. Wasserstein distance learning in hyperbolic spaces builds upon adapting the Wasserstein distance [Villani, 2009] to the structure of hyperbolic geometry [Lee, 2006]. This adaptation involves redefining the cost function in the optimization problem to incorporate the geodesic distance within the hyperbolic manifold. The resulting Wasserstein metric maintains its intuitive interpretation as a measure of the “effort” required to transform one distribution into another while under the manifold constraints of hyperbolic structure. It has proven valuable in various fields, for example, in shape analysis [Shi et al., 2016] and hierarchy matching [Alvarez-Melis et al., 2020, Hoyos-Idrobo, 2020]. Recently, the work in Bonet et al. [2023] proposed extending the sliced-Wasserstein distance [Rabin et al., 2012] to hyperbolic representation, enabling efficient computation of the Wasserstein distance. Our work, however, diverges fundamentally from these approaches. We focus on observational data that inherently possess a hierarchy rather than probability distributions defined on hyperbolic spaces. Additionally, our focus is on the TWD, aiming for efficient computation while addressing the specific challenges posed by the hidden feature hierarchy.
Existing TWDs. The TWD is a popular method to reduce the computational complexity associated with calculating the Wasserstein distance on the Euclidean metric. This approach approximates the original Euclidean metric using a tree metric, as described in several studies [Indyk and Thaper, 2003, Le et al., 2019, Evans and Matsen, 2012, Leeb, 2018, Takezawa et al., 2021, Yamada et al., 2022]. The resulting methods provide a coarser approximation of OT in an Euclidean space, facilitating more efficient computations from cubic complexity to linear complexity. One notable method is the Quadtree [Indyk and Thaper, 2003], which creates a random tree by recursively dividing hypercubes. This tree is then used to compute the TWD. Flowtree [Backurs et al., 2020] modifies the Quadtree method by focusing on the optimal flow and evaluating the cost of this flow within the ground metric. The Tree-Sliced Wasserstein Distance (TSWD) [Le et al., 2019] represents another TWD technique. It computes the average of TWDs by randomly sampling trees. The number of sampling trees is a hyperparameter, commonly set to one, five, or ten, resulting in the variants TSWD-1, TSWD-5, and TSWD-10. Further advancements are presented in Yamada et al. [2022], which introduce WQTWD and WCTWD. These methods employ a Quadtree or clustering-based tree to approximate the Wasserstein distances by optimizing the weights on the tree. Lastly, UltraTree [Chen et al., 2024] constructs an ultrametric tree [Johnson, 1967] by minimizing the OT regression cost, aiming to approximate the Wasserstein distance within the original metric space.
@@
Tree Construction by Graph-Based and HC Methods. Another line of related work involves constructing trees using existing graph-based algorithms, such as Minimum Spanning Tree (MST) [Alon et al., 1995], or hierarchical clustering (HC) methods, particularly those that leverage hyperbolic geometry. Generic graph-based algorithms, like those used in existing TWDs, aim to approximate the original metric with a tree metric. In the context of hierarchical clustering, various methods have been proposed. UltraFit [Chierchia and Perret, 2019] addresses an “ultrametric fitting” problem using Euclidean embeddings. gHHC [Monath et al., 2019] operates under the assumption that partial information about the optimal clustering, specifically the hyperbolic embeddings of the leaves, is known. HHC [Chami et al., 2020] decodes a tree from a two-dimensional hyperbolic embedding for hierarchical clustering by optimizing the continuous Dasgupta’s cost [Dasgupta, 2016]. While these methods may offer theoretical guarantees regarding clustering quality, they do not provide a theoretical guarantee for recovering the latent hierarchy underlying high-dimensional features, as our method does. Our approach uniquely focuses on revealing the latent feature hierarchy, ensuring a more accurate representation of the latent hierarchical structure underlying high-dimensional features.
Fréchet Mean on Manifolds. The Fréchet mean [Fréchet, 1948] is a powerful tool for data analysis on Riemannian manifolds [Lee, 2006], including hyperbolic, spherical, and various matrix manifolds. Its application extends across a diverse range of fields, including batch normalization [Brooks et al., 2019, Lou et al., 2020], batch effect removal [Lin et al., 2021], domain adaptation [Yair et al., 2019], shape analysis and object recognition [McLeod et al., 2012], to name but a few. In our work, we follow this line of work, using the Fréchet mean for constructing the HD-LCA in a product manifold of hyperbolic spaces (see Def. 4.1 and Def. 4.3). Then, this HD-LCA serves as a primary building block in our binary tree decoding (Alg. 1), which in turn, is used in the computation of the proposed TWD (Alg. 2).
@@ Appendix C Theoretical Analysis and Proofs
The hyperbolic LCA $\mathbf{z}_{j}^{k}\vee\mathbf{z}_{j^{\prime}}^{k}$ in Def. 4.1 has a closed-form solution, given by
$\mathbf{z}_{j}^{k}\vee\mathbf{z}_{j^{\prime}}^{k}=\left[\frac{1}{2}\left(\bm{\psi}_{j}^{k}+\bm{\psi}_{j^{\prime}}^{k}\right)^{\top},\mathtt{proj}(\mathbf{z}_{j}^{k}\vee\mathbf{z}_{j^{\prime}}^{k})\right]^{\top},$
where $\mathtt{proj}(\mathbf{z}_{j}^{k}\vee\mathbf{z}_{j^{\prime}}^{k})=\left\lVert\left[\frac{1}{2}\left(\bm{\psi}_{j}^{k}-\bm{\psi}_{j^{\prime}}^{k}\right)^{\top},2^{k/2-2}\right]^{\top}\right\rVert_{2}$.
@@
The Riemannian mean defined using the Fréchet mean [Fréchet, 1948] has a closed-form expression, and is located at the midpoint of the geodesic curve
@@
By Prop. 4.1, we have
$\mathbf{z}_{j}^{k_{1}}\vee\mathbf{z}_{j^{\prime}}^{k_{1}}=\left[\frac{1}{2}\left(\bm{\psi}_{j}^{k_{1}}+\bm{\psi}_{j^{\prime}}^{k_{1}}\right)^{\top},\mathtt{proj}(\mathbf{z}_{j}^{k_{2}}\vee\mathbf{z}_{j^{\prime}}^{k_{1}})\right]^{\top}$
and
$\mathbf{z}_{j}^{k_{2}}\vee\mathbf{z}_{j^{\prime}}^{k_{2}}=\left[\frac{1}{2}\left(\bm{\psi}_{j}^{k_{2}}+\bm{\psi}_{j^{\prime}}^{k_{2}}\right)^{\top},\mathtt{proj}(\mathbf{z}_{j}^{k_{2}}\vee\mathbf{z}_{j^{\prime}}^{k_{2}})\right]^{\top},$
respectively. At the $k_{1}$-th level, the Riemannian distance between the $k_{1}$-th hyperbolic LCA $\mathbf{z}_{j}^{k_{1}}\vee\mathbf{z}_{j^{\prime}}^{k_{1}}$ and $\mathbf{z}_{j}^{k_{1}}$ at the $k_{1}$-th level is given by
@@
The orthogonal projections $\{\mathbf{o}^{k}_{j,j^{\prime}}\}_{k=0}^{K_{c}}$ lie on the unique geodesic path $\gamma_{0}$ that is on the $(m+1)$-axis in $\mathbb{H}^{m+1}$. Note that the geodesic distance joining any two points $\mathbf{o}^{k_{1}}_{j,j^{\prime}}$ and $\mathbf{o}^{k_{2}}_{j,j^{\prime}}$ is given by $\left\lvert\ln\left(\frac{\mathtt{proj}(\mathbf{z}_{j}^{k_{1}}\vee\mathbf{z}_{j^{\prime}}^{k_{1}})}{\mathtt{proj}(\mathbf{z}_{j}^{k_{2}}\vee\mathbf{z}_{j^{\prime}}^{k_{2}})}\right)\right\rvert$ [Stahl, 1993, Udriste, 2013]. The Fréchet mean [Fréchet, 1948] of $\{\mathbf{o}^{k}_{j,j^{\prime}}\}_{k=0}^{K_{c}}$ is the point that minimizes the sum of squared distances to all points in the set. For our set of points, the Fréchet mean mean will have the form $\left[0,\ldots,0,a_{j,j^{\prime}}\right]^{\top}$, where $a_{j,j^{\prime}}\in\mathbb{R}_{>0}$. We want to find $a_{j,j^{\prime}}$ that minimizes
@@
Let $r$ be the root node of $B$ obtained in Alg. 1 and $\mathbf{z}_{r}$ be the corresponding embedding in $\mathcal{M}$. The HD-LCA depth is equivalent to the Gromov product [Gromov, 1987], i.e., $d_{\mathcal{M}}(\mathbf{z}_{r},\overline{\mathbf{h}}_{j,j^{\prime}})\simeq\langle j,j^{\prime}\rangle_{r}$, where $\langle j,j^{\prime}\rangle_{r}=\frac{1}{2}(d_{B}(j,r)+d_{B}(j^{\prime},r)-d_{B}(j,j^{\prime}))$.
@@
Let $\mathbf{z}_{j}$ and $\mathbf{z}_{j^{\prime}}$ be two points in $\mathcal{M}$. We denote $\gamma_{\mathbf{z}_{j}\leadsto\mathbf{z}_{j^{\prime}}}(t)$ the unique geodesic path from $\mathbf{z}_{j}$ to $\mathbf{z}_{j^{\prime}}$ such that $\gamma_{\mathbf{z}_{j}\leadsto\mathbf{z}_{j^{\prime}}}(0)=\mathbf{z}_{j}$ and $\gamma_{\mathbf{z}_{j}\leadsto\mathbf{z}_{j^{\prime}}}(1)=\mathbf{z}_{j^{\prime}}$. Following Bowditch [2007], for any point $\mathbf{z}_{j^{\prime\prime}}\in\mathcal{M}$, we have $d_{\mathcal{M}}(\mathbf{z}_{j^{\prime\prime}},\mathbf{z})\geq\langle j,j^{\prime}\rangle_{j^{\prime\prime}}$, where $\mathbf{z}\in\gamma_{\mathbf{z}_{j}\leadsto\mathbf{z}_{j^{\prime}}}(t)$. Note that if $\mathbf{z}_{j^{\prime\prime}}\in\gamma_{\mathbf{z}_{j}\leadsto\mathbf{z}_{j^{\prime}}}(t)$, then $\langle j,j^{\prime}\rangle_{j^{\prime\prime}}=0$. In addition, let $\gamma_{\mathbf{z}_{j}\leadsto\mathbf{z}_{j^{\prime\prime}}}(t)$ and $\gamma_{\mathbf{z}_{j^{\prime}}\leadsto\mathbf{z}_{j^{\prime\prime}}}(t)$ be the unique geodesic path from $\mathbf{z}_{j}$ to $\mathbf{z}_{j^{\prime\prime}}$ and the unique geodesic path from $\mathbf{z}_{j^{\prime}}$ to $\mathbf{z}_{j^{\prime\prime}}$, respectively. Let $\triangle_{j,j^{\prime},j^{\prime\prime}}=(\gamma_{\mathbf{z}_{j}\leadsto\mathbf{z}_{j^{\prime}}}(t),\gamma_{\mathbf{z}_{j}\leadsto\mathbf{z}_{j^{\prime\prime}}}(t),\gamma_{\mathbf{z}_{j^{\prime}}\leadsto\mathbf{z}_{j^{\prime\prime}}}(t))$ be the triangle connected by these geodesic paths. As $\mathcal{M}$ is 00-hyperbolic, the triangle $\triangle_{j,j^{\prime},j^{\prime\prime}}$ is 0-centre and by triangle inequality, we have
@@
We adopt the equivalence proof from Lin et al. [2023], which stems from Coifman and Lafon [2006] and [Leeb and Coifman, 2016]. Note that diffusion operator $\mathbf{P}$ is typically used for manifold learning, where it be viewed as a discrete proxy of the underlying manifold. It is shown that the diffusion operator recovers the underlying manifold [Coifman and Lafon, 2006], i.e., $\mathbf{P}\overset{m\rightarrow\infty,\epsilon\rightarrow 0}{\longrightarrow}\mathbf{H}=\exp(-\Delta)$, where $\mathbf{H}$ is the Neumann heat kernel of the underlying manifold and $\Delta$ is the Laplace–Beltrami operator on the manifold. In addition, the diffusion operator $\mathbf{P}$ is based on a finite set of points, which is a discrete approximation of $\mathbf{H}$.
@@
Note that for small enough $\epsilon>0$, we have $d_{\mathbf{P},\mathcal{M}}(j,j^{\prime})\simeq d_{\mathbf{H},\mathcal{M}}(j,j^{\prime})$. Specifically, following Coifman and Lafon [2006], we have $\left\lVert\mathbf{P}-\mathbf{H}\right\rVert_{L^{2}(\mathcal{M})}\rightarrow 0$ and $\left\lVert\mathbf{P}-\mathbf{H}\right\rVert_{L^{1}(\mathcal{M})}\rightarrow 0$ as $\epsilon\rightarrow 0$. We denote $d_{\mathcal{M}}(j,j^{\prime})=d_{\mathbf{P},\mathcal{M}}(j,j^{\prime})\simeq d_{\mathbf{H},\mathcal{M}}(j,j^{\prime})$ for abbreviation.
@@
Let $\mu_{i}$ and $\mu_{i^{\prime}}$ be two probability distributions supported on $T$ with a ground metric $d_{T}$, $\Pi(\mu_{i},\mu_{i^{\prime}})$ be the set of couplings $\pi$ on the product space $T\times T$ such that $\pi(B_{1}\times T)=\mu_{i}(B_{1})$ and $\pi(T\times B_{2})=\mu_{i^{\prime}}(B_{2})$ for any sets $B_{1},B_{2}\subset T$. The Wasserstein distance [Villani, 2009] between $\mu_{i}$ and $\mu_{i^{\prime}}$ is defined by
@@
Let $\mathcal{F}_{d_{T}}$ and $\left\lVert\cdot\right\rVert_{L_{d_{T}}}$ denote the Lipschitz functions w.r.t. $d_{T}$ and Lipschitz norm w.r.t. $d_{T}$, respectively. The Kantorovich-Rubinstein dual [Villani, 2009] of Eq. (11) is given by
@@
By Evans and Matsen [2012], the witness function $f$ can be represented by a Borel function $g:T\rightarrow[-1,1]$, i.e.,
@@
where $\lambda$ is the unique Borel measure on $T$. Therefore, following Le et al. [2019], the Wasserstein distance $\mathtt{OT}(\mu_{i},\mu_{i^{\prime}},d_{T})$ can be written as
@@
Note that based on Thm. 4.1, the decoded tree metric $d_{B}(\cdot,\cdot)$ and the underlying tree metric $d_{T}(\cdot,\cdot)$ are bilipschitz equivalent. Additionally, the TWD using the tree $T$ is the Wasserstein distance with ground pairwise distance $d_{T}$ [Evans and Matsen, 2012], i.e., $\mathtt{TW}(\mu_{1},\mu_{2},T)=\mathtt{OT}(\mu_{1},\mu_{2},d_{T})$. Therefore, for any two discrete histograms $\mathbf{x}_{i}$ and $\mathbf{x}_{i^{\prime}}$ supported on $T$, we have $\mathtt{TW}(\mathbf{x}_{i},\mathbf{x}_{i^{\prime}},B)\simeq\mathtt{TW}(\mathbf{x}_{i},\mathbf{x}_{i^{\prime}},T)$. ∎

### v3

@@ A.4 Preliminaries on Trees
Definition A.8 (LCA Relation (Wang & Wang, 2018).).
@@ Appendix B Additional Related Work
Wasserstein Distance on Hyperbolic Manifold. Wasserstein distance learning in hyperbolic spaces builds upon adapting the Wasserstein distance (Villani, 2009) to the structure of hyperbolic geometry (Lee, 2006). This adaptation involves redefining the cost function in the optimization problem to incorporate the geodesic distance within the hyperbolic manifold. The resulting Wasserstein metric maintains its intuitive interpretation as a measure of the “effort” required to transform one distribution into another while under the manifold constraints of hyperbolic structure. It has proven valuable in various fields, for example, in shape analysis (Shi et al., 2016) and hierarchy matching (Alvarez-Melis et al., 2020; Hoyos-Idrobo, 2020). Recently, the work in Bonet et al. (2023) proposed extending the sliced-Wasserstein distance (Rabin et al., 2012) to hyperbolic representation, enabling efficient computation of the Wasserstein distance. Our work, however, diverges fundamentally from these approaches. We focus on observational data that inherently possess a hierarchy rather than probability distributions defined on hyperbolic spaces. Additionally, our focus is on the TWD, aiming for efficient computation while addressing the specific challenges posed by the hidden feature hierarchy.
Existing TWDs. The TWD is a popular method to reduce the computational complexity associated with calculating the Wasserstein distance on the Euclidean metric. This approach approximates the original Euclidean metric using a tree metric, as described in several studies (Indyk & Thaper, 2003; Le et al., 2019; Evans & Matsen, 2012; Leeb, 2018; Takezawa et al., 2021; Yamada et al., 2022). The resulting methods provide a coarser approximation of OT in an Euclidean space, facilitating more efficient computations from cubic complexity to linear complexity. One notable method is the Quadtree (Indyk & Thaper, 2003), which creates a random tree by recursively dividing hypercubes. This tree is then used to compute the TWD. Flowtree (Backurs et al., 2020) modifies the Quadtree method by focusing on the optimal flow and evaluating the cost of this flow within the ground metric. The Tree-Sliced Wasserstein Distance (TSWD) (Le et al., 2019) represents another TWD technique. It computes the average of TWDs by randomly sampling trees. The number of sampling trees is a hyperparameter, commonly set to one, five, or ten, resulting in the variants TSWD-1, TSWD-5, and TSWD-10. Further advancements are presented in Yamada et al. (2022), which introduce WQTWD and WCTWD. These methods employ a Quadtree or clustering-based tree to approximate the Wasserstein distances by optimizing the weights on the tree. Lastly, UltraTree (Chen et al., 2024) constructs an ultrametric tree (Johnson, 1967) by minimizing the OT regression cost, aiming to approximate the Wasserstein distance within the original metric space.
@@
Tree Construction by Graph-Based and HC Methods. Another line of related work involves constructing trees using existing graph-based algorithms, such as Minimum Spanning Tree (MST) (Alon et al., 1995), or hierarchical clustering (HC) methods, particularly those that leverage hyperbolic geometry. Generic graph-based algorithms, like those used in existing TWDs, aim to approximate the original metric with a tree metric. In the context of hierarchical clustering, various methods have been proposed. UltraFit (Chierchia & Perret, 2019) addresses an “ultrametric fitting” problem using Euclidean embeddings. gHHC (Monath et al., 2019) operates under the assumption that partial information about the optimal clustering, specifically the hyperbolic embeddings of the leaves, is known. HHC (Chami et al., 2020) decodes a tree from a two-dimensional hyperbolic embedding for hierarchical clustering by optimizing the continuous Dasgupta’s cost (Dasgupta, 2016). While these methods may offer theoretical guarantees regarding clustering quality, they do not provide a theoretical guarantee for recovering the latent hierarchy underlying high-dimensional features, as our method does. Our approach uniquely focuses on revealing the latent feature hierarchy, ensuring a more accurate representation of the latent hierarchical structure underlying high-dimensional features.
Fréchet Mean on Manifolds. The Fréchet mean (Fréchet, 1948) is a powerful tool for data analysis on Riemannian manifolds (Lee, 2006), including hyperbolic, spherical, and various matrix manifolds. Its application extends across a diverse range of fields, including batch normalization (Brooks et al., 2019; Lou et al., 2020), batch effect removal (Lin et al., 2021; 2024), domain adaptation (Yair et al., 2019), shape analysis and object recognition (McLeod et al., 2012), to name but a few. In our work, we follow this line of work, using the Fréchet mean for constructing the HD-LCA in a product manifold of hyperbolic spaces (see Def. 4.1 and Def. 4.3). Then, this HD-LCA serves as a primary building block in our binary tree decoding (Alg. 1), which in turn, is used in the computation of the proposed TWD (Alg. 2).
@@ Appendix C Theoretical Analysis and Proofs
The hyperbolic LCA $\mathbf{z}_{j}^{k}\vee\mathbf{z}_{j^{\prime}}^{k}$ in Def. 4.1 has a closed-form solution, given by $\mathbf{z}_{j}^{k}\vee\mathbf{z}_{j^{\prime}}^{k}=\left[\frac{1}{2}\left(\bm{\psi}_{j}^{k}+\bm{\psi}_{j^{\prime}}^{k}\right)^{\top},\mathtt{proj}(\mathbf{z}_{j}^{k}\vee\mathbf{z}_{j^{\prime}}^{k})\right]^{\top}$, where $\mathtt{proj}(\mathbf{z}_{j}^{k}\vee\mathbf{z}_{j^{\prime}}^{k})=\left\lVert\left[\frac{1}{2}\left(\bm{\psi}_{j}^{k}-\bm{\psi}_{j^{\prime}}^{k}\right)^{\top},2^{k/2-2}\right]^{\top}\right\rVert_{2}$.
@@
The Riemannian mean defined using the Fréchet mean (Fréchet, 1948) has a closed-form expression, and is located at the midpoint of the geodesic curve
@@
By Prop. 4.1, we have $\mathbf{z}_{j}^{k_{1}}\vee\mathbf{z}_{j^{\prime}}^{k_{1}}=\left[\frac{1}{2}\left(\bm{\psi}_{j}^{k_{1}}+\bm{\psi}_{j^{\prime}}^{k_{1}}\right)^{\top},\mathtt{proj}(\mathbf{z}_{j}^{k_{2}}\vee\mathbf{z}_{j^{\prime}}^{k_{1}})\right]^{\top}$ and $\mathbf{z}_{j}^{k_{2}}\vee\mathbf{z}_{j^{\prime}}^{k_{2}}=\left[\frac{1}{2}\left(\bm{\psi}_{j}^{k_{2}}+\bm{\psi}_{j^{\prime}}^{k_{2}}\right)^{\top},\mathtt{proj}(\mathbf{z}_{j}^{k_{2}}\vee\mathbf{z}_{j^{\prime}}^{k_{2}})\right]^{\top}$, respectively. At the $k_{1}$-th level, the Riemannian distance between the $k_{1}$-th hyperbolic LCA $\mathbf{z}_{j}^{k_{1}}\vee\mathbf{z}_{j^{\prime}}^{k_{1}}$ and $\mathbf{z}_{j}^{k_{1}}$ at the $k_{1}$-th level is given by
@@
The orthogonal projections $\{\mathbf{o}^{k}_{j,j^{\prime}}\}_{k=0}^{K_{c}}$ lie on the unique geodesic path $\gamma_{0}$ that is on the $(m+1)$-axis in $\mathbb{H}^{m+1}$. Note that the geodesic distance joining any two points $\mathbf{o}^{k_{1}}_{j,j^{\prime}}$ and $\mathbf{o}^{k_{2}}_{j,j^{\prime}}$ is given by $\left\lvert\ln\left(\frac{\mathtt{proj}(\mathbf{z}_{j}^{k_{1}}\vee\mathbf{z}_{j^{\prime}}^{k_{1}})}{\mathtt{proj}(\mathbf{z}_{j}^{k_{2}}\vee\mathbf{z}_{j^{\prime}}^{k_{2}})}\right)\right\rvert$ (Stahl, 1993; Udriste, 2013). The Fréchet mean (Fréchet, 1948) of $\{\mathbf{o}^{k}_{j,j^{\prime}}\}_{k=0}^{K_{c}}$ is the point that minimizes the sum of squared distances to all points in the set. For our set of points, the Fréchet mean mean will have the form $\left[0,\ldots,0,a_{j,j^{\prime}}\right]^{\top}$, where $a_{j,j^{\prime}}\in\mathbb{R}_{>0}$. We want to find $a_{j,j^{\prime}}$ that minimizes
@@
Let $r$ be the root node of $B$ obtained in Alg. 1 and $\mathbf{z}_{r}$ be the corresponding embedding in $\mathcal{M}$. The HD-LCA depth is equivalent to the Gromov product (Gromov, 1987), i.e., $d_{\mathcal{M}}(\mathbf{z}_{r},\overline{\mathbf{h}}_{j,j^{\prime}})\simeq\langle j,j^{\prime}\rangle_{r}$, where $\langle j,j^{\prime}\rangle_{r}=\frac{1}{2}(d_{B}(j,r)+d_{B}(j^{\prime},r)-d_{B}(j,j^{\prime}))$.
@@
Let $\mathbf{z}_{j}$ and $\mathbf{z}_{j^{\prime}}$ be two points in $\mathcal{M}$. We denote $\gamma_{\mathbf{z}_{j}\leadsto\mathbf{z}_{j^{\prime}}}(t)$ the unique geodesic path from $\mathbf{z}_{j}$ to $\mathbf{z}_{j^{\prime}}$ such that $\gamma_{\mathbf{z}_{j}\leadsto\mathbf{z}_{j^{\prime}}}(0)=\mathbf{z}_{j}$ and $\gamma_{\mathbf{z}_{j}\leadsto\mathbf{z}_{j^{\prime}}}(1)=\mathbf{z}_{j^{\prime}}$. Following Bowditch (2007), for any point $\mathbf{z}_{j^{\prime\prime}}\in\mathcal{M}$, we have $d_{\mathcal{M}}(\mathbf{z}_{j^{\prime\prime}},\mathbf{z})\geq\langle j,j^{\prime}\rangle_{j^{\prime\prime}}$, where $\mathbf{z}\in\gamma_{\mathbf{z}_{j}\leadsto\mathbf{z}_{j^{\prime}}}(t)$. Note that if $\mathbf{z}_{j^{\prime\prime}}\in\gamma_{\mathbf{z}_{j}\leadsto\mathbf{z}_{j^{\prime}}}(t)$, then $\langle j,j^{\prime}\rangle_{j^{\prime\prime}}=0$. In addition, let $\gamma_{\mathbf{z}_{j}\leadsto\mathbf{z}_{j^{\prime\prime}}}(t)$ and $\gamma_{\mathbf{z}_{j^{\prime}}\leadsto\mathbf{z}_{j^{\prime\prime}}}(t)$ be the unique geodesic path from $\mathbf{z}_{j}$ to $\mathbf{z}_{j^{\prime\prime}}$ and the unique geodesic path from $\mathbf{z}_{j^{\prime}}$ to $\mathbf{z}_{j^{\prime\prime}}$, respectively. Let $\triangle_{j,j^{\prime},j^{\prime\prime}}=(\gamma_{\mathbf{z}_{j}\leadsto\mathbf{z}_{j^{\prime}}}(t),\gamma_{\mathbf{z}_{j}\leadsto\mathbf{z}_{j^{\prime\prime}}}(t),\gamma_{\mathbf{z}_{j^{\prime}}\leadsto\mathbf{z}_{j^{\prime\prime}}}(t))$ be the triangle connected by these geodesic paths. As $\mathcal{M}$ is 00-hyperbolic, the triangle $\triangle_{j,j^{\prime},j^{\prime\prime}}$ is 0-centre and by triangle inequality, we have
@@
We adopt the equivalence proof from Lin et al. (2023), which stems from Coifman & Lafon (2006) and (Leeb & Coifman, 2016). Note that diffusion operator $\mathbf{P}$ is typically used for manifold learning, where it be viewed as a discrete proxy of the underlying manifold. It is shown that the diffusion operator recovers the underlying manifold (Coifman & Lafon, 2006), i.e., $\mathbf{P}\overset{m\rightarrow\infty,\epsilon\rightarrow 0}{\longrightarrow}\mathbf{H}=\exp(-\Delta)$, where $\mathbf{H}$ is the Neumann heat kernel of the underlying manifold and $\Delta$ is the Laplace–Beltrami operator on the manifold. In addition, the diffusion operator $\mathbf{P}$ is based on a finite set of points, which is a discrete approximation of $\mathbf{H}$.
@@
Note that for small enough $\epsilon>0$, we have $d_{\mathbf{P},\mathcal{M}}(j,j^{\prime})\simeq d_{\mathbf{H},\mathcal{M}}(j,j^{\prime})$. Specifically, following Coifman & Lafon (2006), we have $\left\lVert\mathbf{P}-\mathbf{H}\right\rVert_{L^{2}(\mathcal{M})}\rightarrow 0$ and $\left\lVert\mathbf{P}-\mathbf{H}\right\rVert_{L^{1}(\mathcal{M})}\rightarrow 0$ as $\epsilon\rightarrow 0$. We denote $d_{\mathcal{M}}(j,j^{\prime})=d_{\mathbf{P},\mathcal{M}}(j,j^{\prime})\simeq d_{\mathbf{H},\mathcal{M}}(j,j^{\prime})$ for abbreviation.
@@
Let $\mu_{i}$ and $\mu_{i^{\prime}}$ be two probability distributions supported on $T$ with a ground metric $d_{T}$, $\Pi(\mu_{i},\mu_{i^{\prime}})$ be the set of couplings $\pi$ on the product space $T\times T$ such that $\pi(B_{1}\times T)=\mu_{i}(B_{1})$ and $\pi(T\times B_{2})=\mu_{i^{\prime}}(B_{2})$ for any sets $B_{1},B_{2}\subset T$. The Wasserstein distance (Villani, 2009) between $\mu_{i}$ and $\mu_{i^{\prime}}$ is defined by
@@
Let $\mathcal{F}_{d_{T}}$ and $\left\lVert\cdot\right\rVert_{L_{d_{T}}}$ denote the Lipschitz functions w.r.t. $d_{T}$ and Lipschitz norm w.r.t. $d_{T}$, respectively. The Kantorovich-Rubinstein dual (Villani, 2009) of Eq. (11) is given by
@@
By Evans & Matsen (2012), the witness function $f$ can be represented by a Borel function $g:T\rightarrow[-1,1]$, i.e.,
@@
where $\lambda$ is the unique Borel measure on $T$. Therefore, following Le et al. (2019), the Wasserstein distance $\mathtt{OT}(\mu_{i},\mu_{i^{\prime}},d_{T})$ can be written as
@@
Note that based on Thm. 4.1, the decoded tree metric $d_{B}(\cdot,\cdot)$ and the underlying tree metric $d_{T}(\cdot,\cdot)$ are bilipschitz equivalent. Additionally, the TWD using the tree $T$ is the Wasserstein distance with ground pairwise distance $d_{T}$ (Evans & Matsen, 2012), i.e., $\mathtt{TW}(\mu_{1},\mu_{2},T)=\mathtt{OT}(\mu_{1},\mu_{2},d_{T})$. Therefore, for any two discrete histograms $\mathbf{x}_{i}$ and $\mathbf{x}_{i^{\prime}}$ supported on $T$, we have $\mathtt{TW}(\mathbf{x}_{i},\mathbf{x}_{i^{\prime}},B)\simeq\mathtt{TW}(\mathbf{x}_{i},\mathbf{x}_{i^{\prime}},T)$. ∎